a random sample of 100 babies is obtained and the mean head circumference is found to be 4.06cm.Assuming that the population standard deviation is known to be 1.6cm, use 0.05 significance level to test the claim that the mean head circumference of all two-month old babies is equal to 40.0cm
Answer by stanbon(75887)

A random sample of 100 babies is obtained and the mean head circumference is found to be 4.06cm.
-----------------------------
Assuming that the population standard deviation is known to be 1.6cm, use 0.05 significance level to test the claim that the mean head circumference of all two-month old babies is equal to 40.0cm
------------
Ho: u = 4.0 cm
Ha: u is not equal to 4.0 cm
-----------------
t(4.0) = (4.06-4)/[1.6/sqrt(100)] = 0.06/0.16 = 0.375
----------------
p-value = 2*P(0.375< t < 100, when df = 99) = 2*0.3542 = 0.7085
---
Conclusion: Since the p-value is greater tan 5%, fail to reject Ho.
=========================
Cheers,
Stan H.